# Beyond Homophily: Community Search on Heterophilic Graphs

### v1

@@ Abstract
Community search aims to identify a refined set of nodes that are most relevant to a given query, supporting tasks ranging from fraud detection to recommendation. Unlike homophilic graphs, many real-world networks are heterophilic, where edges predominantly connect dissimilar nodes. Therefore, structural signals that once reflected smooth, low-frequency similarity now appear as sharp, high-frequency contrasts. However, both classical algorithms (e.g., $k$-core, $k$-truss) and recent ML-based models struggle to achieve effective community search on heterophilic graphs, where edge signs or semantics are generally unknown. Algorithm-based methods often return communities with mixed class labels, while GNNs, built on homophily, smooth away meaningful signals and blur community boundaries. Therefore, we propose Adaptive Community Search (AdaptCS), a unified framework featuring three key designs: (i) an AdaptCS Encoder that disentangles multi-hop and multi-frequency signals, enabling the model to capture both smooth (homophilic) and contrastive (heterophilic) relations; (ii) a memory-efficient low-rank optimization that removes the main computational bottleneck and ensures model scalability; and (iii) an Adaptive Community Score (ACS) that guides online search by balancing embedding similarity and topological relations. Extensive experiments on both heterophilic and homophilic benchmarks demonstrate that AdaptCS outperforms the best-performing baseline by an average of 11% in F1-score, retains robustness across heterophily levels, and achieves up to 2 orders of magnitude speedup.
@@ I Introduction
Graphs serve as a crucial representation for complex relational data across various domains, such as social networks [tan2023higher, coclep], citation networks [ics-gnn, qdgnn], and molecular structures [molecular_22_nips, rong2020self]. Identifying a closely interrelated community based on a query node has been an important research topic within the database domain. Existing community search (CS) methods can be broadly categorized into algorithm-based and machine learning (ML)-based approaches. Algorithm-based approaches define communities based on structural cohesiveness, such as $k$-core [k-core, fang2020survey], $k$-truss [k-truss, k-truss2017], and $k$-clique [k-clique, k-clique2018], identifying densely interconnected nodes through graph-theoretic measures and optimization criteria. In contrast, ML-based approaches are task-driven and leverage predictive models, explicitly defining communities using known node labels or types [ics-gnn, qdgnn, coclep]. By harnessing learned embeddings, ML-based methods effectively identify nodes related to a query node, emphasizing semantic similarity and class consistency.
Traditional graph algorithms and neural models typically assume homophily, where connected nodes are likely to share similar attributes or belong to the same community [graphsage, gin_2018, gat_2017]. However, real-world graphs often violate this assumption and exhibit heterophily, where edges predominantly link nodes with different labels or communities [heter_freq_filter_aaai_23, heter_luan2023graph, heter_unsupervised_aaai_23]. For instance, in citation networks, nodes represent papers and edges denote citations: while within-domain citations reflect homophily, cross-domain ones are common and crucial for knowledge transfer. For example, a biology paper may cite database works on subgraph matching to analyze protein structures, and medical papers often reference AI vision models for radiology image analysis. Similar heterophilic patterns occur in many real scenarios: fraudsters tend to interact with legitimate users rather than other fraudsters [fraud_jianke_2023, linBP_2015_VLDB], political discussions frequently occur across opposing viewpoints [h2gcn_2020beyond, li2023signed], and proteins in molecular graphs consist of diverse amino acids with distinct properties [h2gcn_2020beyond, li2023homogcl].
@@
Existing solutions. Although no prior work has directly addressed CS under heterophily, both the database and AI communities have extensively studied heterophilic node classification tasks. Building upon these insights, existing methods relevant to heterophilic CS can be broadly categorized into three paradigms: (a) Algorithm-based approaches that rely purely on structural cohesiveness (e.g., $k$-core, $k$-truss, $k$-clique) and are not heterophily-aware; (b) Compatibility-matrix-based label propagation methods [linBP_2015_VLDB, factorLP_2020_SIGMOD], originally developed for heterophilic node classification, model cross-class relations through a compatibility matrix. We extend this line of work to the CS by using the inferred compatibility patterns to guide community expansion around the query. (c) ML-based general solutions, where heterophily-oriented GNN frameworks (e.g., FAGCN [FAGCN21], ACM [ACM_2022], ALT [ALT_KDD_2023]), originally proposed for node classification, are integrated as extensions into ML-based CS models to handle heterophily. These paradigms provide valuable insights yet exhibit fundamental limitations when applied to heterophilic CS.
@@
(b) Compatibility-matrix based label propagation ($H$-based). This line of methods [linBP_2015_VLDB, factorLP_2020_SIGMOD], tackles heterophily in node classification by introducing a pre-defined or statistically learned compatibility matrix that governs how labels interact across edges. The inferred matrix enables cross-class propagation beyond the homophily assumption and can be naturally extended to community search by expanding from a query node to others predicted to share the same community. However, the compatibility matrix $H$ is either predefined or globally optimized, remaining fixed across the graph and unable to adapt to local variations in heterophily. As illustrated in Figure 1(b), the query node $v_{q}$ simultaneously connects to two distinct neighborhoods, where one is dominated by positive links and the other by negative links, exhibiting contrasting local homophily levels that a single global $H$ cannot accurately capture. Consequently, such models often mispropagate signals across incompatible regions and fail to retrieve communities with heterophilic edge semantics.
(c) ML-based CS (heterophilic extension). Recent ML-based community search models (e.g., ICSGNN [ics-gnn], QDGNN [qdgnn], COCLEP [coclep]) are developed under an implicit homophily assumption and thus fail to generalize to heterophilic graphs. Although general heterophily-aware GNNs such as FAGCN, ACM, and ALT can be applied as extensions to these models, their aggregation remains distance-agnostic: multi-hop signals are recursively blended across layers, giving rise to what we define as the Flip Effect (III.1) in multi-class settings. As illustrated in Figure 1(c), consider nodes $v_{1}$ and $v_{2}$ that are both heterophilic neighbors of the query node $v_{q}$. Because the model mixes messages across hops, the two-step path $v_{1}\overset{-}{\longleftrightarrow}v_{q}\overset{-}{\longleftrightarrow}v_{2}$ introduces a false positive relation between $v_{1}$ and $v_{2}$ despite their different labels (node colors).
@@ II Related Work
Algorithm-based community search. Traditional community search algorithms utilize various cohesiveness metrics to identify communities within graphs. Metrics such as $k$-core [k-core, tan2023higher], $k$-truss [k-truss, k-truss2017], and $k$-clique [k-clique, k-clique2018] have been employed to efficiently detect subgraphs that meet predefined structural criteria. For instance, the $k$-core metric identifies subgraphs where each node has at least $k$ connections within the subgraph, ensuring a level of internal connectivity. $k$-truss focuses on the presence of triangles, identifying subgraphs where each edge participates in at least k2k-2 triangles, thus capturing a higher-order cohesiveness. These methods have been extended to attributed graphs, incorporating node attributes alongside structural considerations to identify communities of nodes sharing similar characteristics [attributed, attributed1].
@@
Machine learning-based community search. The advent of GNNs has introduced flexible and expressive models for community search, capable of balancing contributions from both topological structures and node attributes. Models like ICSGNN [ics-gnn] leverage GNNs to capture similarities between nodes by combining content and structural features. This approach allows for the interactive and iterative discovery of target communities, guided by user feedback. Similarly, QDGNN [qdgnn] employs an offline setting, training on a fixed dataset and extending to attributed community search by adopting an attribute encoder to identify groups of nodes with specific attributes. Other models, such as ALICE [jianwei_sig] and COCLEP [coclep], incorporate advanced techniques like cross-attention encoders and contrastive learning to enhance the expressiveness and efficiency of community search in attributed graphs. More recently, SMN [smn_25] and CommunityDF [chen2025communitydf] further advance this line of research: SMN proposes a general solution for overlapping community search via subspace embedding, while CommunityDF introduces a generative diffusion-based framework that iteratively refines query-centered subgraphs through contrastive learning and dynamic thresholding.
@@
Learning on heterophilic graphs. Another research line extends classical label propagation to heterophilic settings through a learnable compatibility matrix that models class interactions across edges. LinBP [linBP_2015_VLDB] reformulates belief propagation into a linear system supporting both homophilic and heterophilic relations via a global matrix $H$, while FactorLP [factorLP_2020_SIGMOD] further learns $H$ from labeled data through matrix factorization. These approaches preserve the interpretability and efficiency of propagation but remain globally linear, unable to adapt to local heterophily or leverage node attributes.
@@
Beyond propagation models, recent GNN-based methods aim to enhance robustness under heterophily. Approaches such as Geom-GCN [geom-gcn], MixHop [Mixhop], and GPRGNN [GRPGNN_ICLR_2021] expand or reweight message passing to capture high-order and flexible dependencies, while FAGCN [FAGCN21] introduces signed edge weighting to model cross-class relations. General frameworks like ACM [ACM_2022] and ALT [ALT_KDD_2023] further extend these ideas by adaptively combining multi-channel propagation or optimizing graph structure, enabling GNNs to handle graph heterophily.
@@ III Preliminaries
Let G=(V,E)G=({V,E}) be an undirected graph with a set V{V} of nodes and a set E{E} of edges. Let n=|V|n=|{V}| and m=|E|m=|{E}| be the number of nodes and edges, respectively. Given a node $u\in{V}$, ${N}_{u}=\{v|(u,v)\in{E}\}$ is the neighbor set of $u$. The adjacency matrix of $G$ is denoted as ${A}\in\{0,1\}^{n\times n}$, where ${A}_{i,j}=1\text{, if }(v_{i},v_{j})\in{E}$, otherwise ${A}_{i,j}=0$. ${X}\in\mathbb{R}^{n\times d}$ is the set of node features, where $d$ is the dimension of the feature, and ${x}_{i}$ represents the node features of $v_{i}$. We use ${Z}\in\mathbb{R}^{n\times c}$ to denote the label encoding matrix, whose $i$-th row is the one-hot encoding of the label of $v_{i}$ and $c$ is the dimension of the label. Given a query node $q$, the CS problem aims to find a $\mathcal{K}$-sized set of nodes containing the query from $G$ while maximizing the embedding similarity learned by ML models against the query [ics-gnn, qdgnn, coclep].
@@ III-A Problem Definition
A heterophilic graph exhibits a low $h_{edge}(G)$, meaning that most edges connect nodes with different community labels. Following the existing definition of the CS in ML-based models [ics-gnn, qdgnn, coclep, smn_25], we define the CS as below:
@@ III-B Graph Laplacian
The graph Laplacian is defined as L=DAL=D-A, which is symmetric and positive semidefinite. Its eigen-decomposition can be expressed as $L=U\,\Lambda\,U^{\top}$, where $U\in\mathbb{R}^{n\times n}$ is a matrix whose columns are orthonormal eigenvectors (forming the graph Fourier basis) and $\Lambda=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{n})$ is a diagonal matrix holding the corresponding eigenvalues (or frequencies) arranged such that $0\leq\lambda_{1}\leq\cdots\leq\lambda_{n}$. Variations include the degree normalized Laplacian (a symmetric matrix), defined as $\tilde{L}_{\mathrm{sym}}=D^{-1/2}\,L\,D^{-1/2}=I-D^{-1/2}{A}D^{-1/2}$, and random walk normalized Laplacian (non-symmetric), defined as $\tilde{L}_{\mathrm{rw}}=D^{-1}\,L\,=I-D^{-1}{A}$. In graph signal processing, the Laplacian $L$ and its variants serve as high-pass filters, while the corresponding affinity (or adjacency) matrices, such as $\tilde{A}_{\mathrm{sym}}=I-\tilde{L}_{\mathrm{sym}}=D^{-1/2}AD^{-1/2}$, act as low-pass filters. A renormalization trick is to add self-loops by defining $\bar{A}=A+I$ and $\bar{D}=D+I$; the renormalized matrices are then given by $\hat{A}_{\mathrm{sym}}=\bar{D}^{-1/2}\bar{A}\bar{D}^{-1/2}$ and $\hat{L}_{\mathrm{sym}}=I-\hat{A}_{\mathrm{sym}}$. In GCN framework [gcn2016], the output is computed as
@@
where $W_{0}\in\mathbb{R}^{f\times f_{1}}$ and $W_{1}\in\mathbb{R}^{f_{1}\times o}$ are learnable weight matrices. The random walk renormalized matrix $\hat{A}_{\mathrm{rw}}=\bar{D}^{-1}\bar{A}$ and its corresponding Laplacian $\hat{L}_{\mathrm{rw}}=I-\hat{A}_{\mathrm{rw}}$ serve as mean aggregators in spatial-based GNNs, sharing same eigenvalues as $\hat{A}_{\mathrm{sym}}$. We use $\hat{A}_{\mathrm{sym}}$ in this paper and denote it as $\hat{A}$.
@@ III-C Flip Effect
Given a labeled graph $G=(V,E)$, let $z_{u}$ denote the community label of node $u\in V$. Suppose the graph is heterophilic, such that edges $(u,v)\in E$ often connect nodes with different labels ($z_{u}\neq z_{v}$). The flip effect arises when aggregating information across multiple hops: if there exists a two-hop path $u\overset{-}{\longleftrightarrow}w\overset{-}{\longleftrightarrow}v$ with $z_{u}\neq z_{w}$ and $z_{w}\neq z_{v}$, standard aggregation may falsely infer $z_{u}=z_{v}$, thereby “flipping” the underlying relationship.
@@
Multi-class case. With the four-class case, ideally, the red color should only appear on the diagonal (highlighted regions). However, the same aggregation produces additional off-diagonal red blocks, notably between Class 1 and Class 3, and between Class 2 and Class 4. These correspond to node pairs that are connected via an even number of negative edges (such as two-hop paths), causing their similarity to be incorrectly flipped from negative to positive. As a result, Laplacian-based high-frequency filtering fails to distinguish communities in the multi-class setting, merging structurally distinct groups and demonstrating the ”flip effect”—an inherent limitation of high-pass aggregation in general heterophilic graphs.

### v2

@@ Abstract
Community search aims to identify a refined set of nodes that are most relevant to a given query, supporting tasks ranging from fraud detection to recommendation. Unlike homophilic graphs, many real-world networks are heterophilic, where edges predominantly connect dissimilar nodes. Therefore, structural signals that once reflected smooth, low-frequency similarity now appear as sharp, high-frequency contrasts. However, both classical algorithms (e.g., $k$-core, $k$-truss) and recent ML-based models struggle to achieve effective community search on heterophilic graphs, where edge signs or semantics are generally unknown. Algorithm-based methods often return communities with mixed class labels, while GNNs, built on homophily, smooth away meaningful signals and blur community boundaries. Therefore, we propose Adaptive Community Search (AdaptCS), a lightweight framework featuring three key designs: (i) an AdaptCS Encoder that disentangles multi-hop and multi-frequency signals, enabling the model to capture both smooth (homophilic) and contrastive (heterophilic) relations; (ii) a memory-efficient low-rank optimization that removes the main computational bottleneck and ensures model scalability; and (iii) an Adaptive Community Score (ACS) that guides online search by balancing embedding similarity and topological relations. Extensive experiments on both heterophilic and homophilic benchmarks demonstrate that AdaptCS outperforms the best-performing baseline by an average of 11% in F1-score, retains robustness across heterophily levels, and achieves up to 2 orders of magnitude speedup over the strongest ML-based CS baselines.
@@ I Introduction
Graphs serve as a crucial representation for complex relational data across various domains, such as social networks [1, 2], citation networks [3, 4], and molecular structures [5, 6]. Identifying a closely interrelated community based on a query node has been an important research topic within the database domain. Existing community search (CS) methods can be broadly categorized into algorithm-based and machine learning (ML)-based approaches. Algorithm-based approaches define communities based on structural cohesiveness, such as $k$-core [7, 8], $k$-truss [9, 10], and $k$-clique [11, 12], identifying densely interconnected nodes through graph-theoretic measures and optimization criteria. In contrast, ML-based approaches are task-driven and leverage predictive models, explicitly defining communities using known node labels or types [13, 4, 14]. By harnessing learned embeddings, ML-based methods effectively identify nodes related to a query node, emphasizing semantic similarity and class consistency.
Traditional graph algorithms and neural models typically assume homophily, where connected nodes are likely to share similar attributes or belong to the same community [15, 16, 17]. However, real-world graphs often violate this assumption and exhibit heterophily, where edges predominantly link nodes with different labels or communities [18, 19, 20]. For instance, in citation networks, nodes represent papers and edges denote citations: while within-domain citations reflect homophily, cross-domain ones are common and crucial for knowledge transfer. For example, a biology paper may cite database works on subgraph matching to analyze protein structures, and medical papers often reference AI vision models for radiology image analysis. Similar heterophilic patterns occur in many real scenarios: fraudsters tend to interact with legitimate users rather than other fraudsters [21, 22], political discussions frequently occur across opposing viewpoints [23, 24], and proteins in molecular graphs consist of diverse amino acids with distinct properties [23, 25]. Notably, the edge signs in these scenarios are typically not explicitly observed, which limits the applicability of signed-graph models that assume known positive/negative relations.
@@
Existing solutions. Although no prior work has directly addressed CS under heterophily, both the database and AI communities have studied graph heterophily on node classification tasks. Building upon these insights, existing methods relevant to heterophilic CS can be broadly categorized into three paradigms: (a) Algorithm-based approaches that rely purely on structural cohesiveness (e.g., $k$-core, $k$-truss, $k$-clique) and are not heterophily-aware; (b) Compatibility-matrix-based label propagation methods [22, 26], originally developed for heterophilic node classification, model cross-class relations through a compatibility matrix. We extend this line of work to the CS by using the inferred compatibility patterns to guide community expansion around the query. (c) ML-based general solutions, where heterophily-oriented GNN frameworks (e.g., FAGCN [27], ACM [28], ALT [29]), originally proposed for node classification, are integrated as extensions into ML-based CS models to handle heterophily.
The above paradigms provide valuable insights yet exhibit fundamental limitations when applied to heterophilic CS.
@@
(b) Compatibility-matrix based label propagation ($H$-based). This line of methods [22, 26], tackles heterophily in node classification by using a pre-defined or statistically learned compatibility matrix that governs how labels interact across edges. The inferred matrix supports propagation across different classes without relying on the homophily assumption, and it can be adapted for community search by starting from a query node and expanding to other nodes predicted to have the same community. However, the compatibility matrix $H$ is either predefined or globally optimized, remaining fixed across the graph and unable to adapt to local variations in heterophily. As shown in Figure 1(b), the query node $v_{q}$ connects to two distinct communities, where one is dominated by positive links and the other by negative links, exhibiting contrasting local homophily levels that a single global $H$ cannot accurately capture. Consequently, such models often mispropagate signals across incompatible regions and fail to retrieve communities with heterophilic edge semantics.
(c) ML-based CS (heterophilic extension). Recent ML-based community search models (e.g., ICSGNN [13], QDGNN [4], COCLEP [14]) are developed under an implicit homophily assumption and thus fail to generalize to heterophilic graphs. Although general heterophily-aware GNNs such as FAGCN, ACM, and ALT can be applied as extensions to these models, their aggregation remains distance-agnostic: multi-hop signals are recursively blended across layers, giving rise to what we define as the Flip Effect (III.1) in multi-class settings. As illustrated in Figure 1(c), consider nodes $v_{1}$ and $v_{2}$ that are both heterophilic neighbors of the query node $v_{q}$. Because the model mixes messages across hops, the two-step path $v_{1}\overset{-}{\longleftrightarrow}v_{q}\overset{-}{\longleftrightarrow}v_{2}$ introduces a false positive relation between $v_{1}$ and $v_{2}$ despite their different labels (different node colors).
@@ II Related Work
Algorithm-based community search. Traditional community search algorithms utilize various cohesiveness metrics to identify communities within graphs. Metrics such as $k$-core [7, 1], $k$-truss [9, 10], and $k$-clique [11, 12] have been employed to efficiently detect subgraphs that meet predefined structural criteria. For instance, the $k$-core metric identifies subgraphs where each node has at least $k$ connections within the subgraph, ensuring a level of internal connectivity. $k$-truss focuses on the presence of triangles, identifying subgraphs where each edge participates in at least k2k-2 triangles, thus capturing a higher-order cohesiveness. These methods have been extended to attributed graphs, incorporating node attributes alongside structural considerations to identify communities of nodes sharing similar characteristics [30, 31].
@@
Machine learning-based community search. The advent of GNNs has introduced flexible and expressive models for community search, capable of balancing contributions from both topological structures and node attributes. Models like ICSGNN [13] leverage GNNs to capture similarities between nodes by combining content and structural features. This approach allows for the interactive and iterative discovery of target communities, guided by user feedback. Similarly, QDGNN [4] employs an offline setting, training on a fixed dataset and extending to attributed community search by adopting an attribute encoder to identify groups of nodes with specific attributes. Other models, such as ALICE [32] and COCLEP [14], incorporate advanced techniques like cross-attention encoders and contrastive learning to enhance the expressiveness and efficiency of community search in attributed graphs. More recently, SMN [33] and CommunityDF [34] further advance this line of research: SMN proposes a general solution for overlapping community search via subspace embedding, while CommunityDF introduces a generative diffusion-based framework that iteratively refines query-centered subgraphs through contrastive learning and dynamic thresholding.
@@
Learning on heterophilic graphs. Another research line extends classical label propagation to heterophilic settings through a learnable compatibility matrix that models class interactions across edges. LinBP [22] reformulates belief propagation into a linear system supporting both homophilic and heterophilic relations via a global matrix $H$, while FactorLP [26] further learns $H$ from labeled data through matrix factorization. These approaches preserve the interpretability and efficiency of propagation but remain globally linear, unable to adapt to local heterophily or leverage node attributes.
@@
Beyond propagation models, recent GNN-based methods aim to enhance robustness under heterophily. Approaches such as Geom-GCN [35], MixHop [36], and GPRGNN [37] expand or reweight message passing to capture high-order and flexible dependencies, while FAGCN [27] introduces signed edge weighting to model cross-class relations. General frameworks like ACM [28] and ALT [29] further extend these ideas by adaptively combining multi-channel propagation or optimizing graph structure, enabling GNNs to handle graph heterophily.
@@ III Preliminaries
Let G=(V,E)G=({V,E}) be an undirected graph with a set V{V} of nodes and a set E{E} of edges. Let n=|V|n=|{V}| and m=|E|m=|{E}| be the number of nodes and edges. Given a node $u\in{V}$, ${N}_{u}=\{v|(u,v)\in{E}\}$ is the neighbor set of $u$. The adjacency matrix of $G$ is denoted as ${A}\in\{0,1\}^{n\times n}$, where ${A}_{i,j}=1\text{, if }(v_{i},v_{j})\in{E}$, otherwise ${A}_{i,j}=0$. ${X}\in\mathbb{R}^{n\times d}$ is the set of node features, where $d$ is the dimension of the feature, and ${x}_{i}$ represents the node features of $v_{i}$. We use ${Z}\in\mathbb{R}^{n\times c}$ to denote the label matrix, whose $i$-th row is the one-hot encoding of the label of $v_{i}$ and $c$ is the dimension of the label.
@@ III-A Problem Definition
A heterophilic graph exhibits a low $h_{edge}(G)$, meaning that most edges connect nodes with different community labels. Following the existing definition of the CS in ML-based models [13, 4, 14, 33], we define the CS as below:
@@ III-B Graph Laplacian
In heterophilic graphs, low-pass aggregation with normalized adjacency can over-mix features from different labels. To preserve heterophily-informative signals, our encoder additionally uses a high-pass channel. The graph Laplacian is the standard high-pass operator; thus we briefly review it below and use it to construct our high-frequency propagation.
The graph Laplacian is defined as L=DAL=D-A, where $D$ is the degree matrix, which is symmetric and positive semidefinite. Variations include the degree normalized Laplacian (a symmetric matrix), defined as $\tilde{L}_{\mathrm{sym}}=D^{-1/2}\,L\,D^{-1/2}=I-D^{-1/2}{A}D^{-1/2}$, and random walk normalized Laplacian (non-symmetric), defined as $\tilde{L}_{\mathrm{rw}}=D^{-1}\,L\,=I-D^{-1}{A}$. In GCN framework [38], the low-pass output is computed as
@@
where $W_{0}\in\mathbb{R}^{f\times f_{1}}$ and $W_{1}\in\mathbb{R}^{f_{1}\times o}$ are learnable weight matrices. $\hat{A}_{\mathrm{sym}}$ is the low-pass filter, and can be replaced by the high-pass filter $\hat{L}_{\mathrm{sym}}$ to form a high-frequency channel. The random walk renormalized matrix $\hat{A}_{\mathrm{rw}}=\bar{D}^{-1}\bar{A}$ and its corresponding Laplacian $\hat{L}_{\mathrm{rw}}=I-\hat{A}_{\mathrm{rw}}$ serve as mean aggregators in spatial-based GNNs, sharing same eigenvalues as $\hat{A}_{\mathrm{sym}}$. We use $\hat{A}_{\mathrm{sym}}$ in this paper and denote it as $\hat{A}$.
@@ III-C Flip Effect
Given a multi-label graph $G=(V,E)$, let $z_{u}$ denote the community label of node $u\in V$. Suppose the graph is heterophilic, such that edges often connect nodes with different labels. The flip effect arises when aggregating information across multiple hops: if there exists a two-hop path $u\overset{-}{\longleftrightarrow}w\overset{-}{\longleftrightarrow}v$ with $z_{u}\neq z_{w}$, $z_{w}\neq z_{v}$ and $z_{u}\neq z_{v}$, standard aggregation may falsely infer $z_{u}=z_{v}$, thereby “flipping” the underlying relationship.
@@
Multi-class case. With the four-class case, ideally, the red color should only appear on the diagonal (highlighted inter-class regions). However, the same aggregation produces additional off-diagonal red blocks, notably between Class 1 and Class 3, and between Class 2 and Class 4. These correspond to node pairs that are connected via an even number of negative edges (such as two-hop paths), causing their similarity to be incorrectly flipped from negative to positive. Hence, Laplacian-based high-frequency filtering fails to distinguish communities in the multi-class setting, merging structurally distinct groups and demonstrating the ”flip effect”—an inherent limitation of high-pass aggregation in general heterophilic graphs.